Jim 2nd test score was 8 points higher than his first test score His third score was 88 He had a B average between 80 and 89 inclusive for the three tests What can you conclude about his 1st score? show me how you do it please.
Answer by mathmandotcom(31)

Let x=test1
x+8=test2
88=test3
jim had a b average so 80<=(t1+t2+t3)/3<=89 the avg of his test
substitution gives us
80<=(x+x+8+88)/3<=89; 80<=(2x+96)3<=89; multiply by 3
240<=2x+96<=267; subtract 96
240-96<=2x<=267-96; 144<=2x<=171; now divide by 2
77<=x<=85.5
if jim first test was between 77 and 85.5 inclusive